OpenStudy (anonymous):

solve the equation 2x+1=-5x-8.

OpenStudy (jhannybean):

Move your x-values to one side and your constants to the other. 2x +5x = -8 -1 Smplify both sides 7x = -9 divide by 7 on both sides and oyu will find your answer

OpenStudy (anonymous):